OpenStudy (anonymous):

Can someone solve and show steps for r + s + t = 5 2r - 7s - 3t = 13 1/2 r - 1/3s + 2/3t = -1

OpenStudy (anonymous):

Multiply first equation by 2: 2r+2s+2t=10 Subtract second from first: 2r+2s+2t=10 -(2r-7s-3t)=-(13) -5s-t=-3 Now multiply the third equation by 4: 2r-4/3s+2t=-1 Subtract second from third: 2r-4/3s+2t=-1 -(2r-7s-3t)=-(13) -25/3s-t=-14 Now take your two new equations and subtract: 25/3s-t=-14 -5s-t=-3 40/3s=-11 s=-33/40 Plug in to find s: -5s-t=-3 -5s+33/40=-3 153/40=5s s=153/200 Plug in to find r: 2r+2s+2t=10 2r+2(33/40)+2(153/200)=10 2r=10-318/200 2r=8.41 r=4.205 s=.825 t=.765 I'm pretty sure those numbers are all wrong but this is the idea. I don't have paper to write down and figure this out and its hard to keep track of everything on here, sorry.

OpenStudy (anonymous):

Yea its wrong.